Suppose the real rate is 9 percent and the inflation rate is 1.4 percent. What rate would you expect to see on a Treasury bill?
Answer #1

(1+real rate)=(1+nominal rate)/(1+rate of inflation)

(1+0.09)=(1+nominal rate)/(1+0.014)

1+nominal rate=(1.09*1.014)

nominal rate=(1.09*1.014)-1

=10.526%